A block of mass 30.0 [kg] slides down from the top of the incline. The coefficient of kinetic friction between the block and the incline is πœ‡π‘˜ = 0.350. What is the total work done on the block as it reaches the bottom of the incline?


1
Expert's answer
2022-05-31T10:04:54-0400
"W=mgh-F_f d=mgd\sin\theta-\mu mg d\cos\theta\\
=mgd(\sin\theta-\mu\cos\theta)"